OpenStudy (anonymous):

The theoretical probability of undesirable side effects resulting from taking Grebex is 1 in 11. If 121 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (zarkon):

the answer to your specific question is ...we don't know. if you want the expected number of people that suffer undesirable side effects then use the formula \(n\times p\)

OpenStudy (anonymous):

like this x/121 = 1/11 -- x = 121(1/11) --- x = 11 (# that will encounter unwanted side effects)

OpenStudy (zarkon):

sure...or just \[121\times\frac{1}{11}=11\]

OpenStudy (anonymous):

thank you